OpenStudy (anonymous):

wich of th following is equal to 80+0.1+0.02 85.21 85.12 80.12 80.21

OpenStudy (alexandervonhumboldt2):

80*0.1+0.02=80+0.10+0.02=80+0.12=??????

OpenStudy (anonymous):

80.12

OpenStudy (beccasmith):

80.12